OpenStudy (amy0799):

consider the following function. f(x) = sqrt(9-x^2) Find the derivative from the left at x = 3. Find the derivative from the right at x = 3.

OpenStudy (thadyoung):

find the derivaite, use chain rule, and then plug in 3 for every x

OpenStudy (amy0799):

can you show me how?

OpenStudy (chillout):

Let \(u=sqrt(v)\) and \(v=9-x²\). What does the chain rule tells you?

OpenStudy (chillout):

tell*

OpenStudy (chillout):

\(\frac{d}{dx}f(g(x))=f'(g(x))*g'(x))\)

OpenStudy (castiel):

Make it simpler yourself by writing \[(9-x ^{2})^{\frac{ 1 }{ 2 }}\] Now you could use that derivation rule \[\frac{ d }{ dx }x ^{n}= nx ^{n-1}\] after that you still have to multiply it with derivation of \[\frac{ d }{ dx}(9-x ^{2})\]

OpenStudy (amy0799):

would it be 9-2x?

OpenStudy (castiel):

\[\frac{ d }{ dx }(9-x ^{2})\] =\[\frac{ d }{dx }9-\frac{ d }{dx }x ^{2}\] Now what is a derivative of a constant number like 9 here?

OpenStudy (amy0799):

0

OpenStudy (castiel):

that means \[\frac{ d }{ dx }(9-x ^{2}) = -2x\]

OpenStudy (amy0799):

so then how do i find the derivative from the left and right at x=3?

OpenStudy (jhannybean):

\[\lim_{x\rightarrow 3^{-}} (9-x^2)^{1/2}\]\[\lim_{x\rightarrow 3^+} (9-x^2)^{1/2}\]

OpenStudy (amy0799):

do i just plug in 3 for both equation?

OpenStudy (amy0799):

@ChillOut can you help me

OpenStudy (triciaal):